OpenStudy (anonymous):

The point (3, 2) lies on the line 11x – 6y = 21. Determine if this point is a solution of the system 11x – 6y = 21 9x + 8y = 43

OpenStudy (agreene):

9(3)+8(2)=43 so it lies on the 2nd eqn also, meaning it is a solution to the system.

OpenStudy (anonymous):

i got an aswer of 12 not sure if thats correct

OpenStudy (anonymous):

okay so 9*3=27 and 8*2=16 27+16=43

OpenStudy (agreene):

yes

OpenStudy (anonymous):

okay i didn't

OpenStudy (anonymous):

far enough i did 3*2 at first but its 9 not 2

OpenStudy (anonymous):

thank you